# spacing=nonfrench Sy-FAR: Symmetry-based Fair Adversarial Robustness

### v1

@@ 7 Fairness for Subgroup
Ensuring fairness across all subgroups is a central challenge in ML. As Rothblum and Yona [41] point out, the number of potential subgroups in a classification task grows exponentially with the number of classes . Specifically, the number of possible partitions of classes is given by the Bell number, which already exceeds 100,000 for . Even if we restrict attention to fairness between pairs of disjoint subgroups, the combinatorial explosion remains substantial—there are over 28,000 such pairs when —rendering exhaustive subgroup fairness enforcement computationally infeasible.
To address this, Sy-FAR sidesteps the need to explicitly define or enforce fairness for all possible subgroups. Instead, Sy-FAR leverages a simple yet effective regularizer that promotes pairwise symmetry in the confusion matrix: it directly penalizes asymmetries in the likelihood of misclassifying class as vs. as . Consequently, as we now show, by minimizing these directional imbalances across all class pairs, Sy-FAR implicitly enforces fairness not just at the individual class level, but across any subgroup composed of these classes. This property is powerful: when symmetry holds between all class pairs, it induces symmetry over arbitrary partitions and combinations of these classes. As a result, and as a by product of the regularization, Sy-FAR’s symmetry-based regularization leads to subgroup fairness, without requiring explicit group labels, domain knowledge, or combinatorial enumeration.
@@ 7.1 Defining (Sub)group Symmetry
To define subgroup fairness in classification, we build on the soft confusion matrix defined in §4, where each entry denotes the expected probability of predicting class given true class . Recall that is constructed from softmax probabilities over adversarial examples and each row sums to one. Let be a partition of the classes into disjoint subgroups, with and for . Each subgroup may represent a demographic attribute (e.g., gender), a shared feature (e.g., similar appearance), or any meaningful grouping not necessarily known a priori during training. Using this notation, we now define the notion of symmetry at the (sub)group-level.
@@
where and are the subgroup sizes. This quantity captures the average probability of misclassifying a sample from subgroup into any class in subgroup , adjusted for subgroup sizes.
@@ 7.2 From Individual to Subgroup Symmetry
The Sy-FAR method promotes a novel form of fairness by enforcing symmetry in the confusion matrix, requiring that the probability of misclassifying class as matches the reverse direction, i.e., for all . This criterion, which we refer to as class-level symmetry, is both intuitive and efficient to compute.
To ensure fairness at the group level, however, we must show that the same principle extends to partitions of the label space. Formally, given any partition of the classes, we require that the subgroup misclassification rates are symmetric:
At first glance, enforcing such subgroup symmetry appears to demand evaluating an exponential number of subgroup pairs, which is computationally prohibitive. Remarkably, we prove that our class-level symmetry regularizer is sufficient: enforcing at the class level implies subgroup symmetry fairness across all partitions.
The key insight is that both notions of fairness—class-level and subgroup-level—are derived from the same confusion matrix . Symmetry at the level of individual off-diagonal entries aggregates into symmetric relationships across groups. We now formally state and prove the equivalence between class-level and subgroup-level symmetry.
@@
Conversely, if subgroup fairness holds for all partitions , then must be symmetric.
@@
(1) Class Symmetry Fairness for subgroups. Assume for all , and consider any two distinct subgroups . By definition,
@@
(2) Fairness for all subgroup partitions Class Symmetry. Suppose holds for all partitions. For two classes , examine the following three-part partition:
@@
(Notice that .) By subgroup fairness, , hence . As this equality holds for any pair of classes , is class symmetric. ∎
@@ 7.3 Empirically Assessing Subgroup Fairness
In §6, we empirically showed that Sy-FAR promotes fairness for individuals under evasion attacks: In particular, Sy-FAR increases symmetry as well as source- and target-class fair adversarial robustness. Complementing the theoretical result from §7, we now empirically demonstrate that optimizing for symmetry at the level of individual classes also improves symmetry, and as a by-product, source- and target-group fair adversarial robustness, for subgroups. Notably, these improvements are attained in a manner completely agnostic to the subgroup definitions, as Sy-FAR does not receive any subgroup information as part of its input.
@@
We conduct our evaluation on all three face-recognition setups presented in §5. To study subgroup-level fairness, we focus on two demographic attributes provided in the PubFig dataset [23]: gender (male vs. female) and ethnicity (white vs. non-white). Class-to-subgroup assignments follow directly from the PubFig annotations.
@@
To evaluate robustness and fairness at the level of subgroups, we report robust accuracy for each subgroup separately. The differences between subgroups’ robust accuracies inform us about the degree of unfair source-group robust accuracy. Moreover, because we consider two subgroups in each of the partitions, differences in unfair target-group robust accuracies are similar to the differences in source-group robust accuracy.
@@
Fig. 7 reports subgroup robust accuracy for gender on the PubFig setup between males and females. It can be observed that adversarial training exhibits a small gap between males and females (0.1064), but the overall robust accuracy is low. FAAL and SpecNorm achieve higher overall robust accuracy than adversarial training; however, males attain significantly higher robust accuracy, with the gender gap reaching for FAAL and for SpecNorm. By contrast, Sy-FAR not only achieves the highest overall robust accuracy, but the gap between males and females is the lowest (0.04), with females receiving higher robsut accuracy than males. The results for datasets other than PubFig show identical trends (Table 11 in App. D), albeit Sy-FAR fares best in the challenging PubFig setting. The results for the ethnicity partition (Fig. 12 and Table 12 in App. D) show similar trends as those for the gender partition, with similar levels of robust accuracy across gender and ethnicity groups. Note that on all settings and between all subgroups, the gaps in benign accuracy between subgroups is small (0.04). Overall, these results confirm that symmetry-based regularization not only encourages class-level fair adversarial robustness, but also propagates to subgroup-level fair adversarial robustness, substantially reducing demographic disparities under attacks.
@@ 8 Conclusion
This work introduced Sy-FAR, an approach for enhancing fair source- and target-class adversarial robustness through symmetry. Sy-FAR is based on an the intuitive insight that class resemblance is usually a symmetric relationship, i.e., misclassifications of adversarial examples between classes should be symmetric. We extensively evaluate Sy-FAR, comparing it leading approaches [18, 63], on fairness- and security-critical face-recognition tasks against realistic eyeglass attacks as well as standard object-classification benchmarks commonly used in the literature. Our results, among others, demonstrate substantial improvements in fair source-class adversarial robustness while benign and robust accuracy are preserved or improved; identify unfair target-class robustness as a risk that Sy-FAR helps ameliorate; extend to group-based fairness where we show theoretically and empirically how symmetry advances fairness for arbitrary sub-groups in a computationally efficient manner. Beyond fairness, Sy-FAR offers strong practical advantages: It incurs only negligible overhead relative to standard adversarial training, in sharp contrast to established methods, and it exhibits lower variance across independent runs, highlighting its stability and reproducibility.
@@ Acknowledgments
This work has been supported in part by grant No. 2023641 from the United States-Israel Binational Science Foundation (BSF); by ISF grant no. 1807/23; by Len Blavatnik and the Blavatnik Family foundation; by a Maof prize for outstanding young scientists; by the Ministry of Innovation, Science & Technology, Israel (grant number 0603870071); by the Stellar Development Foundation; and by a grant from the Tel Aviv University Center for AI and Data Science (TAD).
@@ Appendix A Evaluation on CIFAR-10 and CIFAR-100
While our primary focus is on face recognition, where fairness disparities pose immediate security and social risks, we also evaluate Sy-FAR on the widely used CIFAR-10 and CIFAR-100 benchmarks with the PreAct-ResNet18 model. These datasets and model are standard in the adversarial robustness and fairness literature [63, 18], enabling direct comparison with prior methods. Since they lack real-world fairness challenges such as demographic balance or sibling similarity, we present the detailed results here in the Appendix, keeping the main text centered on security-critical face recognition scenarios.
@@ Appendix B Evaluation Under Targeted Eyeglass Attack
Moreover, as summarized in Tables 9 and 10, Sy-FAR strengthens fair source-class adversarial robustness, narrows per-class disparities, and distributes misclassifications more evenly across targets. Overall, these results demonstrate that Sy-FAR preserves robustness while advancing fairness even under targeted attacks.
@@ Appendix C Stability
Figs. 9–11 demonstrate Sy-FAR’s ability to consistently optimize adversarial robustness and fairness across different runs.
@@ Appendix D Additional Subgroup Fairness Results
Complementing the results from §7.3, Fig. 12 shows the robust accuracy for ethnic groups (white vs. non-white) on PubFig, while Table 11 and Table 12 report the benign and robust accuracy gaps between gender and ethnicity groups across all face-recognition setups we consider.

### v2

@@ 7 Fairness for Subgroup
Ensuring fairness across all subgroups is a central challenge in ML. As Rothblum and Yona [rothblum2018probably] point out, the number of potential subgroups in a classification task grows exponentially with the number of classes . Specifically, the number of possible partitions of classes is given by the Bell number, which already exceeds 100,000 for =10. Even if we restrict attention to fairness between pairs of disjoint subgroups, the combinatorial explosion remains substantial—there are over 28,000 such pairs when =10—rendering exhaustive subgroup fairness enforcement computationally infeasible.
To address this, Sy-FAR sidesteps the need to explicitly define or enforce fairness for all possible subgroups. Instead, Sy-FAR leverages a simple yet effective regularizer that promotes pairwise symmetry in the confusion matrix: it directly penalizes asymmetries in the likelihood of misclassifying class as vs. as . Consequently, as we now show, by minimizing directional imbalances across all class pairs, Sy-FAR promotes symmetry not only at the individual class level but also across any subgroup composed of these classes. While symmetry does not explicitly enforce fairness, our experiments demonstrate that encouraging such symmetry consistently leads to improved fairness outcomes at the subgroup level. This property is powerful: when symmetry holds between all class pairs, it induces symmetry over arbitrary partitions and combinations of these classes. As a result, and as a by product of the regularization, Sy-FAR’s symmetry-based regularization enhances subgroup fairness, without requiring explicit group labels, domain knowledge, or combinatorial enumeration.
@@ 7.1 Defining (Sub)group Symmetry
To define subgroup symmetry in classification, we build on the soft confusion matrix defined in §4, where each entry denotes the expected probability of predicting class given true class . Recall that the matrix is constructed from softmax probabilities over adversarial examples and each row sums to one. Let be a partition of the classes into disjoint subgroups, with and for . Each subgroup may represent a demographic attribute (e.g., gender), a shared feature (e.g., similar appearance), or any meaningful grouping not necessarily known a priori during training. Using this notation, we now define the notion of symmetry at the (sub)group-level.
@@
where and are the subgroup sizes.
This quantity captures the average probability of misclassifying a sample from subgroup into any class in subgroup , adjusted for subgroup sizes.
@@ 7.2 From Individual to Subgroup Symmetry
Sy-FAR promotes a novel form of fairness by enforcing symmetry in the confusion matrix, requiring that the probability of misclassifying class as matches the reverse direction, i.e., for all . This criterion, which we refer to as class-level symmetry, is both intuitive and efficient to compute. To ensure fairness at the group level, however, we must show that the same principle extends to partitions of the label space. Formally, given any partition of the classes, we require that any two subgroups satisfy subgroup symmetry:
At first glance, enforcing such subgroup symmetry appears to demand evaluating an exponential number of subgroup pairs, which is computationally prohibitive. Remarkably, we prove that our class-level symmetry regularizer is sufficient: enforcing at the class level implies subgroup symmetry fairness across all partitions and vice versa. The key insight is that both class-level and subgroup-level symmetry arise from the same confusion matrix. Symmetry in individual off-diagonal entries naturally aggregates into symmetric relationships between groups. We now formally state and prove the equivalence between class- and subgroup-level symmetry.
@@
Conversely, if subgroup symmetry holds for all partitions , then must be symmetric.
@@
(1) Class symmetry Symmetry for subgroups.
Assume for all , and consider any two distinct subgroups . By definition,
@@
(2) Symmetry for all subgroup partitions Class Symmetry. Suppose holds for all partitions. For two classes , examine the following three-part partition:
@@
(Notice that .) By subgroup symmetry, , hence . As this equality holds for any pair of classes , is class symmetric. ∎
@@ 7.3 Empirically Assessing Subgroup Fairness
In §6, we empirically showed that Sy-FAR promotes fairness for individuals under evasion attacks; in particular, Sy-FAR increases symmetry as well as source- and target-class fair adversarial robustness. Complementing the theoretical result from §7.2, we now empirically demonstrate that optimizing for symmetry at the level of individual classes also improves symmetry, and as a by-product, source- and target-group fair adversarial robustness, for subgroups. Notably, these improvements are attained in a manner completely agnostic to the subgroup definitions, as Sy-FAR does not receive any subgroup information as part of its input.
@@
We evaluated all three face-recognition setups presented in §5. To study subgroup-level fairness, we focus on two demographic attributes provided in the PubFig dataset [attribute_classifiers]: gender (male vs. female) and ethnicity (white vs. non-white). Class-to-subgroup assignments follow directly from the PubFig annotations.
@@
To evaluate robustness and fairness at the level of subgroups, we report robust accuracy for each subgroup separately. The differences between subgroups’ robust accuracies inform us about the degree of unfair source-group robust accuracy. Note that, because we consider two subgroups in each of the partitions, differences in unfair target-group robust accuracies are similar to the differences in source-group robust accuracy.
@@
Fig. 7 reports subgroup robust accuracy for gender on the PubFig setup between males and females. It can be observed that adversarial training exhibits a small gap between males and females (0.08), but the overall robust accuracy is low. FAAL and SpecNorm achieve higher overall robust accuracy than adversarial training; however, males attain significantly higher robust accuracy, with the gender gap reaching for FAAL and for SpecNorm. By contrast, Sy-FAR not only achieves the highest overall robust accuracy, but the gap between males and females is the lowest (0.04), with females receiving higher robust accuracy than males. The results for datasets other than PubFig show identical trends (Table 11 in App. D), albeit Sy-FAR fares best in the challenging PubFig setting. The results for the ethnicity partition (Fig. 12 and Table 12 in App. D) show similar trends as those for the gender partition, with similar levels of robust accuracy across gender and ethnicity groups. Note that on all settings and between all subgroups, the gaps in benign accuracy between subgroups is small (0.04). Overall, these results confirm that symmetry-based regularization not only encourages class-level fair adversarial robustness, but also propagates to subgroup-level fair adversarial robustness, substantially reducing demographic disparities under attacks.
8 Limitations and Future Work
While Sy-FAR achieves strong and consistent improvements in fair adversarial robustness, several limitations remain, opening promising directions for future research. First, our experiments were conducted under controlled and balanced conditions to isolate the core fairness–robustness challenge. We show that this problem is already difficult in balanced datasets and becomes considerably harder when bias is introduced, as seen in the PubFig setup. Extending Sy-FAR to highly imbalanced or noisy datasets, where labels may be uncertain or ambiguous, is an important next step toward deployment in real-world systems.
Second, the theoretical proof that class-level symmetry implies subgroup-level fairness assumes a finite and correctly labeled class space. In open-world or continuous-class settings—where new or unseen classes may appear—these guarantees may not strictly hold. Future work should explore how to generalize the concept of symmetry to continuous or open-set domains, possibly through embedding-based or probabilistic formulations.
Third, similar to prior work, this work focuses on classification and vision tasks, further extending them to security-critical applications such as face recognition. We believe that the underlying idea of symmetry can also extend to other modalities (e.g., natural language or audio). However, these domains may require domain-specific definitions of symmetric behavior, as similarity definitions and misclassification patterns may differ fundamentally from those in vision.
Finally, extending Sy-FAR to generative and multimodal models will likely require new formalizations of symmetry and fairness that account for continuous output spaces. Investigating these directions constitutes an interesting avenue for future work.
@@ 9 Conclusion
This work introduced Sy-FAR, an approach for enhancing fair source- and target-class adversarial robustness through symmetry. Sy-FAR is based on the intuitive insight that class resemblance is usually a symmetric relationship, i.e., misclassifications of adversarial examples between classes should be symmetric. We extensively evaluate Sy-FAR, comparing it leading approaches [jin2025enhancing, zhang2024towards], on fairness- and security-critical face-recognition tasks against realistic eyeglass attacks as well as standard object-classification benchmarks commonly used in the literature. Our results, among others, demonstrate substantial improvements in fair source-class adversarial robustness while benign and robust accuracy are preserved or improved; identify unfair target-class robustness as a risk that Sy-FAR helps ameliorate; extend to group-based fairness where we show theoretically and empirically how symmetry advances fairness for arbitrary sub-groups in a computationally efficient manner. Beyond fairness, Sy-FAR offers strong practical advantages: It incurs only negligible overhead relative to standard adversarial training, in sharp contrast to established methods, and it exhibits lower variance across independent runs, highlighting its stability and reproducibility.
@@
Ethical Considerations
Our work proposes Sy-FAR, a training method aimed at improving fairness and adversarial robustness of ML models, with a particular focus on face-recognition tasks. Below, we discuss the relevant stakeholders, the ethical implications of our research process and publication, and the measures we took to mitigate potential negative impacts.
Stakeholders.
The stakeholders relevant to this work include:
individuals whose images appear in publicly available datasets used in our experiments;
end users of ML-based security systems that may adopt fairness- and robustness-enhancing methods;
organizations and practitioners developing, deploying, or auditing such systems;
members of society at large, particularly populations that may be disproportionately affected by face recognition or automated decision-making; and
the research community, including researchers building on our methods or findings.
Research Process and Experimental Scope.
We evaluate face- and object-recognition models using only existing, publicly available datasets. Importantly, no human subjects or deployed systems were involved in our experiments. We did not collect new biometric data, scrape online platforms, or attempt to identify or track individuals beyond the labels already present in the datasets. We used each public dataset strictly in accordance with its published usage guidelines and terms, without augmenting it with any additional personal information.
All experiments are conducted offline, and we apply only known adversarial perturbation techniques introduced in prior work. We do not introduce new attack vectors that could increase risk to existing systems. As a result, our work does not expose new vulnerabilities or create additional attack capabilities beyond what is already documented in the adversarial ML literature. We do not experiment with live systems and do not violate any terms of service.
Positive and Negative Potential Impacts.
On the positive side, Sy-FAR has the potential to improve fairness and promote more equal treatment across demographic groups in ML-based security systems. By encouraging more symmetric misclassification behavior and reducing worst-class performance gaps, our work can help identify and mitigate biases that disproportionately harm certain individuals or groups. Our analysis also contributes to academic understanding of fairness–robustness trade-offs, supporting the design and evaluation of fairer and more trustworthy ML systems.
At the same time, we acknowledge that improvements in robustness and fairness may have dual-use implications. More effective face-recognition systems could be misused in surveillance or monitoring applications that threaten privacy or civil liberties, particularly when deployed without transparency or regulation. More generally, technical improvements alone do not address broader social or political concerns regarding the appropriateness of such systems. Our goal is not to promote or legitimize such deployments. Rather, we aim to highlight (un)fairness and robustness issues in existing models and to inform the research community about potential trade-offs and mitigation strategies. We do not endorse the deployment of face-recognition or surveillance systems, particularly in high-risk or sensitive contexts.
Mitigations and Ethical Choices.
To mitigate potential negative impacts, we make several deliberate choices in how we design and conduct our study. We restricted experiments to offline evaluations on public datasets, avoided live systems and real users, and used only known adversarial methods. We did not scrape data, violate terms of service, or introduce new attack techniques. Throughout the paper, we present Sy-FAR primarily as a mitigation approach for these issues during training, rather than as a justification for deploying face-recognition systems. These limitations and caveats are clearly stated to ensure responsible interpretation and use of our results.
Deception and Researcher Wellbeing.
There were no user studies, no deception, and no involvement of outside participants. Consequently, there was no risk of harm to participants. The research did not involve exposure to disturbing content or conditions affecting researcher wellbeing, and standard academic safeguards were sufficient.
Reflection.
We believe it is important to publish this work despite potential dual-use risks, as failing to study and report fairness, robustness issues, and possible mitigations may allow biased or unevenly robust systems to persist unexamined. However, we emphasize that technical solutions alone cannot resolve all ethical challenges surrounding biometric technologies. We hope this work informs both technical research and broader discussions on governance, regulation, and responsible use.
Open Science
Our code repository333https://zenodo.org/records/17901662 or https://github.com/haneenn24/Sy-FAR-Symmetry-based-Fair-Adversarial-Robustness. includes all necessary artifacts to fully reproduce the results of the paper. A detailed README file provides step-by-step instructions on how to set up the environment, run training, and reproduce both our baseline adversarial training and the proposed fairness-aware methods (FAAL [zhang2024towards], SpecNorm [jin2025enhancing], and Sy-FAR). The instructions cover both clean training and adversarial training pipelines, as well as evaluation against benign and adversarial inputs (e.g., created with the eyeglass attack). To facilitate reproduction of results without training models from scratch, in addition to the source code, we provide pre-trained model checkpoints under reported settings trained with different robustness–fairness objectives. All artifacts are documented to allow end-to-end reproduction of experiments, ensuring transparency and verifiability of our contributions.
@@ Acknowledgments
This work has been supported in part by grant No. 2023641 from the United States-Israel Binational Science Foundation (BSF); by ISF grant no. 1807/23; by Len Blavatnik and the Blavatnik Family foundation; by a Maof prize for outstanding young scientists; by the Ministry of Energy, Israel (grant number 224-11-026); by the Ministry of Innovation, Science & Technology, Israel (grant number 1001702114); by the Stellar Development Foundation; and by a grant from the Tel Aviv University Center for AI and Data Science (TAD).
@@ Appendix A Evaluation on CIFAR-10 and CIFAR-100
While our primary focus is on face recognition, where fairness disparities pose immediate security and social risks, we also evaluate Sy-FAR on the widely used CIFAR-10 and CIFAR-100 benchmarks with the PreAct-ResNet18 model. These datasets and model are standard in the adversarial robustness and fairness literature [zhang2024towards, jin2025enhancing], enabling direct comparison with prior methods. Since they lack real-world fairness challenges such as demographic balance or sibling similarity, we present the detailed results here in the Appendix, keeping the main text centered on security-critical face recognition scenarios.
@@ Appendix B Evaluation Under Targeted Eyeglass Attack
Moreover, as summarized in Tables 9–10, Sy-FAR strengthens fair source-class adversarial robustness, narrows per-class disparities, and distributes misclassifications more evenly across targets. Overall, these results demonstrate that Sy-FAR preserves robustness while advancing fairness even under targeted attacks.
@@ Appendix C Stability
Figs. 9–11 demonstrate Sy-FAR consistently optimize adversarial robustness and fairness across different runs.
@@ Appendix D Additional Subgroup Fairness Results
Complementing the results from §7.3, Fig. 12 shows the robust accuracy for ethnic groups (white vs. non-white) on PubFig, while Tables 11–12 report the benign and robust accuracy gaps between gender and ethnicity groups across all face-recognition setups we consider.
@@
Appendix E Face-Recognition Model Training
The training process of the face-recognition models follows prior work [tong2021facesec]. We begin with a pre-trained backbone (i.e., feature extractor), to which we attach a randomly initialized fully connected layer followed by a softmax activation. Subsequently, training is conducted in an end-to-end manner, meaning that all layers of the network are optimized jointly rather than freezing any subset. To account for stochastic variations due to random initialization, the entire training procedure is repeated ten times, each time starting from a new random seed. This repetition allows us to estimate the stability and consistency of the observed results.
Appendix F Scalability and Run-Time Analysis
To evaluate the computational scalability of Sy-FAR, we analyze the complexity and run-time contribution of the proposed symmetry regularization term across datasets with different numbers of classes. During training, nearly all computation time is spent on the forward and backward propagation through the deep network and on adversarial example generation–these steps involve billions of floating-point operations. In contrast, the symmetry penalty performs only lightweight arithmetic operations over class-pair statistics derived from the model’s existing softmax outputs. The penalty does not require additional forward or backward passes, nor does it alter the adversarial optimization process. Although the number of class pairs grows quadratically with the number of classes (), each symmetry computation is an extremely small GPU tensor operation over already available data. Empirically, we find that:
For 10 classes (e.g., PubFig), the symmetry term accounts for 0.001% of the total training time per batch.
For 100 classes (e.g., CIFAR-100), the cost increases slightly to 0.01–0.05%.
Even for 1,000 classes (e.g., ImageNet-scale settings), the overhead remains 0.5%.
In other words, the computation remains negligible compared to the dominant convolutional and adversarial training operations. As the number of classes increases, the total epoch time rises modestly due to the linear growth of the final fully connected (FC) layer and softmax computations, while the symmetry term–despite its theoretical complexity–still contributes a negligible fraction of the total run time.
Table 13 summarizes the empirical estimates. The epoch durations are approximate and provided only to illustrate proportional scaling, not to claim fixed absolute values.
| #Classes (C) | Pairs () | Share of Training Time | Epoch Time (min, ) | Symmetry Time per Epoch |
| 10 | 45 | s | ||
| 100 | 4,950 | – | – s | |
| 1000 | 499,500 | – | – s |
Appendix G Extended Attack Evaluation: Face Masks
To evaluate Sy-FAR under more realistic physical threats, we introduce a face-mask attack inspired by FACESEC [tong2021facesec]. Unlike eyeglass attacks, which apply pixel-level, small, localized, high-frequency perturbations around the eyes [sharif2016accessorize], the face-mask attack operates at the grid level, occluding the lower face and inducing a large, low-frequency perturbation that removes key shape and landmark information. Despite this stronger and fundamentally different threat model, Sy-FAR consistently achieves the best symmetry loss, source- and target-class robustness, and fairness (Tables 14–16), demonstrating that its symmetry-based regularization remains effective even under severe facial occlusion.
| Method | PubFig | PubFig | PubFig | |||
| Asym. Gap | Sym. Loss | Asym. Gap | Sym. Loss | Asym. Gap | Sym. Loss | |
| Adv. Train | 0.6183 | 2.0856 | 0.4048 | 1.1772 | 0.3602 | 1.4638 |
| FAAL | 0.2117 | 0.7686 | 0.3333 | 1.0288 | 0.3962 | 1.8115 |
| SpecNorm | 0.2475 | 0.8215 | 0.3418 | 1.0426 | 0.3462 | 1.5138 |
| Sy-FAR | 0.1748 | 0.4946 | 0.3117 | 0.9276 | 0.2974 | 1.0920 |
| Method | PubFig | PubFig | PubFig | |||||||||
| Clean | Robust | Min | Gap | Clean | Robust | Min | Gap | Clean | Robust | Min | Gap | |
| Adv. Train | 97.05 | 56.35 | 1.48 | 88.01 | 95.60 | 64.18 | 25.33 | 60.45 | 82.26 | 54.46 | 17.65 | 65.69 |
| FAAL | 98.60 | 79.36 | 57.78 | 39.44 | 93.04 | 68.54 | 26.67 | 58.33 | 81.06 | 57.97 | 20.82 | 69.82 |
| SpecNorm | 98.10 | 77.02 | 55.85 | 38.42 | 95.75 | 72.04 | 28.33 | 60.84 | 82.19 | 56.61 | 26.16 | 63.63 |
| Sy-FAR | 99.09 | 84.76 | 72.88 | 27.12 | 95.76 | 77.05 | 44.12 | 55.88 | 83.87 | 65.06 | 33.89 | 58.96 |
| Method | PubFig | PubFig | PubFig | ||||||
| MinTgt | MaxTgt | Std | MinTgt | MaxTgt | Std | MinTgt | MaxTgt | Std | |
| Adv. Train | 0.0049 | 0.2736 | 0.0831 | 0.0000 | 0.2485 | 0.0882 | 0.0088 | 0.3719 | 0.0900 |
| FAAL | 0.0038 | 0.2913 | 0.0904 | 0.0000 | 0.3432 | 0.0950 | 0.0016 | 0.3683 | 0.0872 |
| SpecNorm | 0.0078 | 0.2435 | 0.0801 | 0.0044 | 0.2595 | 0.0899 | 0.0055 | 0.2461 | 0.0933 |
| Sy-FAR | 0.0105 | 0.1991 | 0.0623 | 0.0062 | 0.2146 | 0.0669 | 0.0115 | 0.1907 | 0.0582 |